Determine for which value (s) of λ the real part of z = 1+λi/1−λi equals zero


1
Expert's answer
2021-07-09T09:44:40-0400

"z=\dfrac{1+\lambda {i}}{1-\lambda{i}}\\
z=\dfrac{(1+\lambda{i})(1+\lambda{i})}{(1-\lambda{i})(1+\lambda{i})}\\
z=\dfrac{1+2\lambda{i}-\lambda^2}{1+\lambda^2}\\
z=\dfrac{1-\lambda^2}{1+\lambda^2}+\dfrac{2\lambda{i}}{1+\lambda^2}\\
\reals(z)=\dfrac{1-\lambda^2}{1+\lambda^2}=0\\
\therefore
1-\lambda^2=0\\
\lambda^2=1\\
\lambda=\pm \sqrt{1}\\
\lambda=\pm1"